what mass of lead (ii) iodide will be formed when 20.0 g of lead (ii) nitrate is added to 33.0g of potassium iodide? which is the limiting reactant?


1
Expert's answer
2022-03-10T10:46:09-0500

lead(II) iodide = PbI2

lead(II) nitrate = Pb(NO3)2

potassium iodide = KI


Solution:

The molar mass of KI is 166.0 g/mol

The molar mass of Pb(NO3)2 is 331.2 g/mol


Calculate the moles of each reactant:

Moles of Pb(NO3)2 = (20.0 g Pb(NO3)2) × (1 mol Pb(NO3)2 / 331.2 g Pb(NO3)2) = 0.0604 mol Pb(NO3)2

Moles of KI = (33.0 g KI) × (1 mol KI / 166.0 g KI) = 0.1988 mol KI


Balanced chemical equation:

Pb(NO3)2 + 2KI → PbI2 + 2KNO3

According to stoichiometry:

1 mol of Pb(NO3)2 reacts with 2 mol of KI

Thus, 0.0604 mol of Pb(NO3)2 reacts with:

(0.0604 mol Pb(NO3)2) × (2 mol KI / 1 mol Pb(NO3)2) = 0.1208 mol KI

However, initially there is 0.1988 mol of KI (according to the task).

Thus, Pb(NO3)2 acts as limiting reactant and KI is excess reactant.



According to stoichiometry:

1 mol of Pb(NO3)2 produces 1 mol of PbI2

Thus, 0.0604 mol of Pb(NO3)2 produces:

(0.0604 mol Pb(NO3)2) × (1 mol PbI2 / 1 mol Pb(NO3)2) = 0.0604 mol PbI2


The molar mass of PbI2 is 461.0 g/mol

Therefore,

Mass of PbI2 = (0.0604 mol PbI2) × (461.0 g PbI2 / 1 mol PbI2) = 27.8444 g PbI2 = 27.8 g PbI2

Mass of PbI2 = 27.8 g


Answer:

The limiting reactant is Pb(NO3)2

27.8 grams of lead(II) iodide (PbI2) will be formed